OpenStudy (anonymous):

Need Help ASAP!!! Solve using substitution: y − 2x − 1 = 0 -4y = 6 − 8x

OpenStudy (anonymous):

(1) y-2x-1=0 (2) -4y=6-8x First put them both in standard form. (1) y=2x+1 (2) y=2x-6/4 We see that the slopes are the same which means the lines are parallel; there is no solution.

OpenStudy (anonymous):

thank u